Using a financial calculator, find the present value of the remaining loan payments on this mortgage.

$132,000 Mortgage with IOU Mortgage Bank 5 years ago, 30 year mortgage @ 6% interest and a monthly payment of $791 (Total 2010 interest paid $7,433)
Answer #1

The problem can be solved using a financial calculator as described below:

- Input I/Y = 0.5 % (=6/12), N = 12 x 25 = 300 (25 years of the mortgage left), PMT = $ 791 and FV = $ 0

- CMPT-> PV

- PV = $ 122768.63